calculate the percent composition by mass of each element in KClO2.
1
Expert's answer
2019-09-23T08:20:00-0400

M(KClO2) = 106.5 g/mol

C%(K) = 39/106.5 = 0,366 = 36.6 %

C%(Cl) = 35.5/106.5 = 0.333 = 33.3%

C%(O) = 32/106.5 = 0.300 = 30%